# A Dipolar Chiral Spin Liquid on the Breathed Kagome Lattice

### 26 Mar 2026 v1

@@ VI Probing CSL order in finite-sized clusters
Starting at large applied field $h_{z}\gtrsim 1$, the ground state of the system falls squarely in the paramagnetic phase, with all the spins aligned with the chosen pattern and exhibiting negligible spin-spin and $\langle\chi_{i}\chi_{j}\rangle$ correlations [Fig. 9(a)]. As the local field is reduced to $0.01\lesssim h_{z}\lesssim 0.2$, the cluster begins developing correlations [Fig. 9(c)], until all $\langle\chi_{i}\chi_{j}\rangle$ become ferromagnetic for $h_{z}\lesssim 7\times 10^{-3}$. This is in good agreement with the infinite cylinder geometry value $h_{z,c}^{\text{cyl}}/J=8.5\times 10^{-3}$ (Fig.7). In this region, even the most distant triangles (i.e. $\langle\chi_{D}\chi_{E}\rangle$) exhibit small but non-zero chiral-chiral correlations, signaling the onset of TRS-breaking order.
@@
Experimental Characterization of the CSL cluster
While we characterized the nature of the CSL phase by its response to an applied magnetic flux, such an experiment is quite challenging in current AMO platforms due to the need to implement artificial gauge fields. Although some strategies have been developed [62, 63, 64], it is important to explore alternative signatures to verify the topological properties of the spin liquid state. In particular, we will focus on three distinct signatures and how they can be probed using individual spin control.
@@
Figure 10: (a) Dynamical response of the cluster to a quenched chemical potential gradient $\mu(y)$ in the vertical direction. Throughout the evolution, the magnetization dynamics follows the chemical potential gradient as expected. However, at later times, we can also discern a transversal motion of the magnetization, in agreement with a non-zero transverse conductivity. (b) To quantify this transversal response we compute the total spin imbalance $\Delta S^{z}$ between the left and right sides of the cluster (red). Crucially, the direction of this motion reverses when either we consider the opposite TRS-breaking ground state (blue), or we reverse the sign of the chemical potential gradient (dashed line). (c) By rescaling the transversal spin current $J_{x}$ by the strength of the gradient $E_{y}$, we observe a scaling collapse of the early-time dynamics confirming we are studying the linear response of the system.
While our observations occur before the system reaches steady state, they capture important features of the system’s equilibrium linear response. First, the strength of the observed spin current, $I_{x}=\frac{d\Delta S^{z}}{dt}$, is proportional to the strength of the applied chemical potential difference, $I_{x}\propto E_{y}$. This is in agreement with the system being in a linear response regime (i.e. a non-zero transverse conductivity), rather than dominated by incoherent processes–i.e. Fermi’s Golden rule. Indeed, in this setting, the presence of low-lying excited states arising from the edge mode should preclude such an analysis. Second, the direction of the transverse spin current is controlled by both the sign of the chemical potential gradient $E_{y}$, and also the sign of the spontaneous TRS-breaking of the cluster $\overline{\langle\chi\rangle}$ [Fig. 10(c)]; flipping either one reverses to transverse motion of the magnetization.
@@
Our observations are summarized in Fig. 11[a-c], focusing on the magnetization dynamics measured with respect to the initial ground state: $\Delta S^{z}_{\triangle}(t)=\langle S^{z}_{\triangle}(t)\rangle_{\text{dyn}}-\langle S^{z}_{\triangle}\rangle_{\text{gs}}$. The magnetization is initially confined to a single triangle, then subsequently propagates clockwise, highlighting the chiral nature of the edge mode. While the edge mode is not exactly contained within the outer layer of triangles, most of the magnetization remains far from the center of the cluster.
@@
Figure 12: (a) In an odd cluster, the additional $1/2$ magnetization is carried by an excitation on top of the CSL ground state. By applying a local field on the center of the cluster (colored sites), we can change the nature of this excitation. When the field is negative, the central triangle will trap a localized semion. When the field is zero, the magnetization lives on the edge, occupying a chiral gapless mode. When the field is positive, a semion of opposite magnetization is trapped, and two edge modes are occupied. (b) The magnetization of the center and edge triangles . At no applied field, there is a net $1/2$ magnetization at the edge triangles, consistent with an anyon on the edge. When increasing the applied field, the edge picks up another net $1/2$ magnetization while the center has a 1/2-1/2 magnetization. At negative applied field, the net magnetization travels from the edge to the center. (c) The average chirality $\bar{\langle\chi\rangle}$ as a function of the applied field. When the field is positive, the system remains in the CSL state by adding another excitation to the edge. When the field is negative, the TRS symmmetry is restored and the system is no longer in the CSL state.
We conclude this section by discussing how a semion might be trapped using local controls. The key insight is that, because the semion is a half-spin excitation, its chemical potential is controllable using local detunings. This offers a path towards controlling the dynamics of the semionic excitations, using the same tools as used when preparing the CSL. More specifically, we focus on the effect of a locally changing potential at the center of our cluster: $H_{\text{trap}}=h_{z}\sum_{i\in\text{center}}S^{z}_{i}$.
@@
Possible Experimental Platforms
@@
Tweezer Arrays
@@
Indeed, current dipolar XY Rydberg experiments already have all the ingredients necessary to adiabatically prepare the CSL discussed, from the Hamiltonian to the use of a secondary optical tweezer array to engineer the necessary local magnetic field patterns and measure multi-basis observables [75]. Furthermore, the dynamical experiments proposed in Section VI to measure topological signatures of the CSL (conductivity and the edge mode) are also quite natural and within the scope of current experiments.
@@
While the complexity of the internal molecular structure makes certain experimental aspects more challenging, there has been rapid development in the preparation and manipulation of ultracold molecules in recent years, making this platform a promising future direction [86].
@@ VII Discussion and Outlook
Looking forward, there are a number of interesting experimental and theoretical directions to explore. On the experimental side, we have already shown how local control allows us to perform dynamical experiments that directly probe the transverse conductivity, chiral edge mode, and localization of semionic excitations. With the ability to trap and control semions, the outstanding goal would be to exchange semions and measure their anyonic statistics. Indeed, the measurement and interferometry of the topologically protected chiral edge mode can be utilized in certain robust quantum state transfer protocols [92]. On the practical side, we note that the adiabatic preparation timescale always scale with the size of the system. Developing alternative preparation methods is crucial for the high-fidelity dynamical characterization of the CSL in large system sizes.
@@
Acknowledgements— We gratefully acknowledge the insights of M. Bintz, E. Demler, J. Hauschild, L. Homeier, R. Mong, L. Pollet, and R. Veressen. FM acknowledges support from the NSF through a grant for ITAMP at Harvard University, as well as from the Netherlands Organisation for Scientific Research (NWO/OCW), as part of Quantum Limits (project number SUMMIT.1.1016). SC acknowledges support from the National Science Foundation Graduate Research Fellowship under Grant No. DGE 2140743.

### 15 Apr 2026 v2

@@ III.2 Preparation
Starting at large applied field $h_{z}\gtrsim 1$, the ground state of the system falls squarely in the paramagnetic phase, with all the spins aligned with the chosen pattern and exhibiting negligible spin-spin and $\langle\chi_{i}\chi_{j}\rangle$ correlations [Fig. 8(c)]. As the local field is reduced to $0.01\lesssim h_{z}\lesssim 0.2$, the cluster begins developing correlations [Fig. 8(d)], until all $\langle\chi_{i}\chi_{j}\rangle$ become ferromagnetic for $h_{z}\lesssim 7\times 10^{-3}$. This is in good agreement with the infinite cylinder geometry value $h_{z,c}^{\text{cyl}}/J=8.5\times 10^{-3}$. In this region, even the most distant triangles (i.e. $\langle\chi_{D}\chi_{E}\rangle$) exhibit small but non-zero chiral-chiral correlations, signaling the onset of TRS-breaking order.
@@
III.3 Experimental Signatures
While we characterized the nature of the CSL phase by its response to an applied magnetic flux, such an experiment is quite challenging in current AMO platforms due to the need to implement artificial gauge fields. Although some strategies have been developed [69, 70, 71], it is important to explore alternative signatures to verify the topological properties of the spin liquid state. In particular, we will focus on four distinct signatures and how they can be probed using individual spin control.
@@ III.3 Experimental Signatures
Figure 9: (a) Dynamical response of the cluster to a quenched chemical potential gradient $\mu(y)$ in the vertical direction. Throughout the evolution, the magnetization dynamics follows the chemical potential gradient as expected. However, at later times, we can also discern a transverse motion of the magnetization, in agreement with a non-zero transverse conductivity. (b) To quantify this transverse response we compute the total spin imbalance $\Delta S^{z}$ between the left and right sides of the cluster (red). Crucially, the direction of this motion reverses when either we consider the opposite TRS-breaking ground state (blue), or we reverse the sign of the chemical potential gradient (dashed line). (c) By rescaling the transverse spin current $J_{x}$ by the strength of the gradient $E_{y}$, we observe a scaling collapse of the early-time dynamics confirming we are studying the linear response of the system.
While our observations occur before the system reaches steady state, they capture important features of the system’s equilibrium linear response. First, the strength of the observed spin current, $I_{x}=\frac{d\Delta S^{z}}{dt}$, is proportional to the strength of the applied chemical potential difference, $I_{x}\propto E_{y}$. This is in agreement with the system being in a linear response regime (i.e. a non-zero transverse conductivity), rather than dominated by incoherent processes–i.e. Fermi’s Golden rule. Indeed, in this setting, the presence of low-lying excited states arising from the edge mode should preclude such an analysis. Second, the direction of the transverse spin current is controlled by both the sign of the chemical potential gradient $E_{y}$, and also the sign of the spontaneous TRS-breaking of the cluster $\overline{\langle\chi\rangle}$ [Fig. 9(c)]; flipping either one reverses the transverse motion of the magnetization.
@@
Our observations are summarized in Fig. 10(a-c), focusing on the magnetization dynamics measured with respect to the initial ground state: $\Delta S^{z}_{\triangle}(t)=\langle S^{z}_{\triangle}(t)\rangle_{\text{dyn}}-\langle S^{z}_{\triangle}\rangle_{\text{gs}}$. The magnetization is initially confined to a single triangle, then subsequently propagates clockwise, highlighting the chiral nature of the edge mode. While the edge mode is not exactly contained within the outer layer of triangles, most of the magnetization remains far from the center of the cluster.
@@
Figure 11: (a) In an odd cluster, the additional $1/2$ magnetization is carried by an excitation on top of the CSL ground state. By applying a local field on the center of the cluster (colored sites), we can change the nature of this excitation. When the field is negative, the central triangle will trap a localized semion. When the field is zero, the magnetization lives on the edge, occupying a chiral gapless mode. When the field is positive, a semion of opposite magnetization is trapped, and two edge modes are occupied. (b) The magnetization of the center and edge triangles. At no applied field, there is a net $1/2$ magnetization at the edge triangles, consistent with an anyon on the edge. When increasing the applied field, the edge picks up another net $1/2$ magnetization while the center has a 1/2-1/2 magnetization. At negative applied field, the net magnetization travels from the edge to the center. (c) The average chirality $\bar{\langle\chi\rangle}$ as a function of the applied field. When the field is very positive, the system remains in the CSL state by adding another excitation to the edge. When the field is very negative, the TRS symmetry is restored and the system is no longer in the CSL state.
We now discuss how a semion might be trapped using local controls. The key insight is that, because the semion is a half-spin excitation, its chemical potential is controllable using local applied fields. This offers a path towards controlling the dynamics of the semionic excitations, using the same tools as used when preparing the CSL. More specifically, we focus on the effect of a locally changing potential at the center of our cluster: $H_{\text{trap}}=h_{z}\sum_{i\in\text{center}}S^{z}_{i}$.
@@
Detecting anyonic flux through Wilson loops
Access to local observables also potentiates the study of more complex correlation functions. This opens up the opportunity to study the Wilson loop associated with the CSL state to probe the number of anyons.
To investigate this operator as probe of the CSL state, we first return to the infinite cylinder states obtained in the Sec. II, where the Wilson loop around the cylinder helps to identify the topological sector. Building on previous work [24], we define a candidate Wilson loop along a path $\ell=(a_{1},b_{1},a_{2},b_{2},\ldots)$ as the product of sequential raising and lowering operators:
$W_{\ell}=\prod_{i=1}^{|\ell|/2}S^{+}_{a_{i}}S^{-}_{b_{i}}\penalty 10000\ .$ (11)
where $W_{0}$ denotes the magnitude of the Wilson loop, and $\alpha$ encodes the phase which determines the flux information. Crucially, we observe that $|\langle W_{\ell}\rangle_{\text{ren}}|\gtrsim 0.3$ for all paths considered, suggesting the system is indeed in the deconfined regime. Moreover, the phase information of $\langle W_{\ell}\rangle_{\text{ren}}$ directly measures the anyonic flux in the system: computing it for the two wavefunctions at $\theta=0$ and $\theta=2\pi$, we get symmetric values, $\langle W_{\ell}\rangle_{\text{ren}}^{\theta=0}=0.81$ and $\langle W_{\ell}\rangle_{\text{ren}}^{\theta=2\pi}=-0.80$, corresponding to a flux of $\alpha^{\theta=0}=0$ and $\alpha^{\theta=2\pi}=\pi$, respectively. Note that while there is a phase ambiguity in the definition of $W_{\ell}$, this does not change the phase difference $\Delta\alpha$ between the two states which differentiates between the two topological sectors.
To translate these observations in into an experimental probe, we must ensure that the Wilson loop remains a valid measure, and that its numerical value is large enough to be measured in experiments. In the case of a finite-sized cluster, $W_{\ell}$ does not determine the topological sector of the state but can count the number of anyonic excitations within it. Computing $W_{\ell}$ around the center unit cell we observe a change of sign between finite cluster with zero and large trapping fields, corroborating the previous interpretation of the trapping of a semion. However, the structure of the $W_{\ell}$ for different regions becomes more complex, suggesting that larger system size calculations are important to better understand the robustness of this observation. Moreover, we find that the non-normalized Wilson loop has a magnitude of $\approx 0.3$ per unit cell, enabling the measurement of the relative phase of $W_{\ell}$ for loops interfacing with 4 unit-cells.
III.4 Experimental Platforms
@@
Tweezer Arrays: Neutral Atoms and Molecules
@@
Indeed, current dipolar XY Rydberg experiments already have all the ingredients necessary to adiabatically prepare the CSL discussed, from the Hamiltonian to the use of a secondary optical tweezer array to engineer the necessary local magnetic field patterns and measure multi-basis observables [82]. Furthermore, the proposed dynamical experiments (Figs. 9 and 10) measure topological signatures of the CSL (conductivity and the edge mode) are also quite natural and within the scope of current experiments.
@@
While the complexity of the internal molecular structure makes certain experimental aspects more challenging, there has been rapid development in the preparation and manipulation of ultracold molecules in recent years, making this platform a promising future direction [92, 96].
@@ IV Discussion and Outlook
Looking forward, there are a number of interesting experimental and theoretical directions to explore. On the experimental side, we have already shown how local control allows us to perform dynamical experiments that directly probe the transverse conductivity, chiral edge mode, and localization of semionic excitations. With the ability to trap and control semions, the outstanding goal would be to exchange semions and measure their anyonic statistics. Indeed, the measurement and interferometry of the topologically protected chiral edge mode can be utilized in certain robust quantum state transfer protocols [99]. On the practical side, we note that the adiabatic preparation timescale always scales with the size of the system. Developing alternative preparation methods is crucial for the high-fidelity dynamical characterization of the CSL in large system sizes.
@@
Acknowledgements— We gratefully acknowledge the insights of M. Bintz, E. Demler, J. Hauschild, L. Homeier, R. Mong, L. Pollet, and R. Verresen. FM acknowledges support from the NSF through a grant for ITAMP at Harvard University, as well as from the Netherlands Organisation for Scientific Research (NWO/OCW), as part of Quantum Limits (project number SUMMIT.1.1016). SC acknowledges support from the National Science Foundation Graduate Research Fellowship under Grant No. DGE 2140743.